OpenStudy (anonymous):

someone one update me on how to simplify \[\sqrt{216}\], is it 3sqrt{24}?

OpenStudy (anonymous):

\[\sqrt{216}\]

OpenStudy (anonymous):

is it \[3\sqrt{24}\] is what i meant to type

OpenStudy (phi):

I would factor the 216 into primes: it is divisible by 2, so 2*108, 2* 2* 54, 2*2*2*27, now 27 is divisible by 3 (sum of its digits is div by 3: 2+7= 9) we get 2*2*2* 3*3*3 find pairs of numbers: (2*2) (3*3) 2*3 pull the pairs out from the square root \[ 2\cdot3 \sqrt{2*\cdot3}\] that simplifies to \[ 6\sqrt{6} \]

OpenStudy (phi):

notice you have 3*sqrt(24) but 24 is 2*2* 6 so you could pull out the 2*2 pair to get 3*2*sqrt(6) or 6 sqrt(6)

OpenStudy (anonymous):

perfection, i totally understand now. I needed help on how to do the paring, but now i see whats happening, thank you!

OpenStudy (anonymous):

anything that doesnt pair, you have to mulitply it and put it under the radical

OpenStudy (phi):

yes.

OpenStudy (anonymous):

thank you so much